OpenStudy (anonymous):

Help please! Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 : Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 : Write an equation, in slope intercept form, to represent this data. Part 3 : How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (campbell_st):

y = 3x

OpenStudy (anonymous):

for the first one?

OpenStudy (campbell_st):

to find the gradient m = (54 - 36)/(18-12) m = 3 then to find the y intercept y = 3x + b use 12 and 36 36 = 3 x 12 + b b = 0 hence y = 3x

OpenStudy (campbell_st):

ordered pairs (12,36) and (18, 54)

OpenStudy (campbell_st):

the slope represents the number of calories burnt per minute

OpenStudy (campbell_st):

part 3 sub x = 35 to find y...

OpenStudy (anonymous):

Thanks you guys :)